OpenStudy (anonymous):

The following table shows the probability of rolling the numbers 1 through 6 on a fair die. X-1,2,3,4,5,6 P(X)-1/6,1/6,1/6,1/6,1/6,1/6 What is the standard deviation of the random variable X, "the number that comes up on the die?" A.14.028 B.2.917 C.3.5 D.3.745 E.1.708

OpenStudy (anonymous):

I think it is E. but im not sure,. im kinda confused.

OpenStudy (anonymous):

Did you find the mean already?

OpenStudy (anonymous):

Nope E was just my guess. i have no iea how to do this.. lol

OpenStudy (anonymous):

Haha okay. Here's your process. 1) Find the mean of X. Since you have 6 equally likely possibilities, just add up the 6 possible outcomes and divide by the number of outcomes.

OpenStudy (anonymous):

Following so far? Tell me what mean value you get for X.

OpenStudy (anonymous):

negative.. how do i set up these possibilities.. i dont understand what i add.

OpenStudy (anonymous):

What are the possible outcomes of a single dice roll?

OpenStudy (anonymous):

is it like 1/6*(1)

OpenStudy (anonymous):

1/6*(2)

OpenStudy (anonymous):

then find the mean of the answers

OpenStudy (anonymous):

The possible outcomes are 1,2,3,4,5,and 6. Sum those, divide by 6. That will tell you the AVERAGE value of a single roll.

OpenStudy (anonymous):

3.5

OpenStudy (anonymous):

ohhh,,, so thats it!

OpenStudy (anonymous):

Right. Doesn't that make sense that 3.5 would be the average value of a single roll?

OpenStudy (anonymous):

I know the mean now. I don't know the standard deviation yet. But it's not too hard to figure it out, now that I know the mean.

OpenStudy (anonymous):

Indeed... I was thinking to hard about it,.

OpenStudy (anonymous):

The answer isn't 3.5, by the way. 3.5 is the mean. They asked for the standard deviation. So we're almost there.

OpenStudy (anonymous):

yeah i figured that.. lol its never just that easy.. is finding the standard deviation as easy.

OpenStudy (anonymous):

Not QUITE as easy. But it's not too bad. 4 basic steps. We have 6 possibilities. For each one, find out (x-3.5)^2 (So for example, 1 is one possibility, so I would do (1-3.5^2) = (-2.5)^2 = 6.25) add all of those things up. Divide that by N, the number of possibilities. Square root that to get standard deviation.

OpenStudy (anonymous):

alright,.. that works.

OpenStudy (anonymous):

You understand each step that I'm asking you to do?

OpenStudy (anonymous):

yea im pretty sure i got it.

OpenStudy (anonymous):

Great =D As you do it, feel free to ask me for clarification.

OpenStudy (anonymous):

alright so the second one would look like (2-3.5)=(-1.5)^2=2.5

OpenStudy (anonymous):

Almost right. One tiny mistake. (-1.5)^2 = 2.25

OpenStudy (anonymous):

ahh yeah that what i meant ;P the third one would =0.25

OpenStudy (anonymous):

Right =)

OpenStudy (anonymous):

so would the 4th one. =0.25

OpenStudy (anonymous):

5th =2.25 6th=6.25

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and then add them all up an divide by 6?

OpenStudy (anonymous):

Yes =)

OpenStudy (anonymous):

=2.91

OpenStudy (anonymous):

outstanding!

OpenStudy (anonymous):

One more step!

OpenStudy (anonymous):

awww man

OpenStudy (anonymous):

(square root)

OpenStudy (anonymous):

1.705....

OpenStudy (anonymous):

So yeah. To recap, I know that whole process seemed long and complicated, but it's not too bad. Step 1: Find the mean Step 2: For each possibility, find the difference from the mean and square that. Step 3: Add up all of the results and divide by the number of possibilities. Step 4: Square root that to get SD.

OpenStudy (anonymous):

Yes =)

OpenStudy (anonymous):

MY GUESS WOULD HAVE BEEN RIGHT!

OpenStudy (anonymous):

Yaaaaay! =D You're so smart.

OpenStudy (anonymous):

I know it! :) i am a good guesser! Thank you for teaching me the steps tho i can only get so far with guessing :P

OpenStudy (anonymous):

My pleasure!